OpenStudy (anonymous):

write and equation of the line through (6,-2) and (-3,1)

OpenStudy (anonymous):

First find out what your m value is. As you probably know, this is also called something like "rise over run". ( y2 - y1 ) / ( x2 - x1 ) y2 = 1 y1 = -2 x2 = -3 x1 = 6 ( 1 - (-2) ) / ( (-3) - 6 ) = ( 3 / -9 ) = -1/3 m = -1/3

OpenStudy (anonymous):

so your equation would be y = -1/3x since that line has 0 for a y-intercept. Normally you would have to find out what the value of b as well.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."